# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 3. $(3)$
(3.39) $\begin{split}&R_{t}^{\varepsilon}:=\text{\rm{e}}^{\int_{0}^{t}\langle\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\rangle-\frac{1}{2}\int_{0}^{t}|\eta_{s}^{\varepsilon}|^{2}\text{\rm{d}}s},\\ &\eta_{s}^{\varepsilon}:=\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})\\ &\qquad\quad\times\big{\{}B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{\ast}\mu,P_{s}V(\mu))\big{\}},\ \ s,t\in[0,T].\end{split}$
Since (B) implies (A), (2.24) holds so that there exists a constant $c>0$ such that
$\sup_{x\in\mathbb{R}^{d},z\in\mathbb{B},t\in[0,T]}|B_{t}(x,P_{s}^{*}\mu_{\varepsilon},z)-B_{t}(x,P_{s}^{*}\mu,z)|\leq c\mathbb{W}_{k}(\mu_{\varepsilon},\mu)\leq c\varepsilon\|\phi\|_{L^{k}(\mu)},\ \ \varepsilon\in[0,1].$
Hence, by (B) we find a constant $c(\mu,\phi)>0$ such that
(3.40) $\begin{split}|\eta_{s}^{\varepsilon}|&\leq K\|P_{s}V(\mu_{\varepsilon})-P_{s}V(\mu)\|_{\mathbb{B}}+c\varepsilon\|\phi\|_{L^{k}(\mu)}\\ &\leq c(\mu,\phi)\varepsilon\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}\big{)},\ \ s\in[0,T],\ \varepsilon\in(0,1].\end{split}$
Then by Girsanov’s theorem,
$W_{t}^{\varepsilon}:=W_{t}-\int_{0}^{t}\eta_{s}^{\varepsilon}\text{\rm{d}}s,\ \ s\in[0,T]$
is a Brownian motion under $\mathbb{Q}:=R_{T}^{\varepsilon}\mathbb{P}$. Reformulate (3.9) with $\gamma=\mu$ as
$\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}$ =$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu,P_{t}V(\mu))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}$
=$=\Big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+B_{t}(X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))\Big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t}^{\varepsilon},$
$\qquad\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\mu_{\varepsilon},\mu_{\varepsilon}}=X_{0}^{\mu_{\varepsilon}}=X_{0}^{\mu}+\varepsilon\phi(X_{0}^{\mu}).$
By the weak uniqueness of (3.9) with $\gamma=\mu_{\varepsilon},$ we get
(3.41) $v_{t}^{\varepsilon,2}:=\frac{P_{t}^{\mu_{\varepsilon}}V(\mu_{\varepsilon})-P_{t}^{\mu}V(\mu_{\varepsilon})}{\varepsilon}=\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu_{\varepsilon},\mu})(R_{t}^{\varepsilon}-1)],\ \ t\in[0,T].$
Combinig this with $[V]_{\alpha}\leq 1,$ and using Lemma 2.5 for $p=k^{\ast}$, we obtain
(3.42) $\begin{split}&\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq\frac{1}{\varepsilon}\mathbb{E}[\alpha(|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|)|R_{t}^{\varepsilon}-1|]\\ &\leq\frac{1}{\varepsilon}\Big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{\mu}-X_{t}^{\mu_{\varepsilon},\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}.\end{split}$
By (3.24) and (3.36), we find a constant $c_{0}(\mu,\phi)>0$ such that
(3.43) $\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\leq c_{0}(\mu,\phi),\ \ t\in[0,T],\varepsilon\in(0,1].$
On the other hand, by (3.40), for any $p\geq 1$ there exists a constant $c(p,\mu,\phi)>0$ such that
(3.44) $\mathbb{E}[|R_{t}^{\varepsilon}-1|^{p}]\leq c(p,\mu,\phi)\varepsilon^{p}\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{p}{2}},\ \ t\in[0,T],\varepsilon\in[0,1].$
Combining this with (3.41) and (3.43), we find a constant $c_{1}(\mu,\phi)>0$ such that
$\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}$ $\leq\big{(}\mathbb{E}[\|V(X_{t}^{\mu_{\varepsilon},\mu})\|_{\mathbb{B}}^{k}]\big{)}^{\frac{1}{k}}\frac{\big{(}\mathbb{E}[|R_{t}^{\varepsilon}-1|^{k^{*}}]\big{)}^{\frac{1}{k^{*}}}}{\varepsilon}$
$\leq c_{1}(\mu,\phi)\bigg{(}\int_{0}^{t}\big{(}1+\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\big{)}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.$
This together with (3.33) and (3.37) yields that for some constant $c_{2}(\mu,\phi)>0$,
$\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq 2\|v_{t}^{\varepsilon,1}\|_{\mathbb{B}}^{2}+2\|v_{t}^{\varepsilon,2}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\|v_{s}^{\varepsilon}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)},\ \ t\in[0,T].$
By Grownwall’s inequality, we find a constant $c_{3}(\mu,\phi)>0$ such that
(3.45) $\begin{split}&\|v_{t}^{\varepsilon}\|_{\mathbb{B}}^{2}\leq c_{2}(\mu,\phi)\bigg{(}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}+\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\text{\rm{e}}^{c_{2}(\mu,\phi)(t-s)}\text{\rm{d}}s\bigg{)}\\ &\leq c_{3}(\mu,\phi)\frac{\alpha(t^{\frac{1}{2}})^{2}}{t},\ \ t\in[0,T],\ \ \varepsilon\in(0,1].\end{split}$
(c) For $|\tilde{v}_{s}^{\varepsilon}|$. Recall that $\Theta_{t}=u_{t}+id$ for $u_{t}=u_{t}^{\lambda_{0},\mu}$. Similarly to (2.16) and (2.21), by Itô’s formula for $\Theta_{t}(X_{t}^{\mu_{\varepsilon}})-\Theta_{t}(X_{t}^{\mu})$, we obtain
(3.46) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}^{\varepsilon}=\frac{1}{\varepsilon}\Big{\{}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu_{\varepsilon}})-\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X_{t}^{\mu})\Big{\}}\text{\rm{d}}W_{t}\\ &+\frac{1}{\varepsilon}\Big{[}\lambda_{0}u_{t}(X_{t}^{\mu_{\varepsilon}})-\lambda_{0}u_{t}(X_{0}^{\mu})+\nabla_{B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu_{\varepsilon},P_{t}V(\mu_{\varepsilon}))-B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))}\Theta_{t}(X_{t}^{\mu_{\varepsilon}})\Big{]}\text{\rm{d}}t\\ &+\frac{1}{\varepsilon}[B_{t}(X_{t}^{\mu_{\varepsilon}},P_{t}^{*}\mu,P_{t}V(\mu))-B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,P_{t}V(\mu))]\text{\rm{d}}t,\\ &\tilde{v}_{0}^{\varepsilon}=\frac{\Theta_{0}(X_{0}^{\mu_{\varepsilon}})-\Theta_{0}(X_{0}^{\mu})}{\varepsilon}.\end{split}$
Similarly to (3.30), by combining this with (B), (2.9), (3.45) and the stochastic Gronwall inequality, for any $j\geq 1$ we find constants $c(j),c_{4}(\mu,\phi)>0$ such that
(3.47) $\begin{split}&\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c(j)|\phi(X_{0}^{\mu})|^{j}+c_{4}(\mu,\phi)\\ &\qquad+c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{j+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{j}{j+1}},\ \ t\in[0,T],\varepsilon\in(0,1].\end{split}$
Letting $j=k$, taking expectation to both sides and noting that
$c(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{\frac{k+1}{k}}\text{\rm{d}}s\bigg{)}^{\frac{k}{k+1}}$
$\leq\tilde{c}(j)\bigg{(}\int_{0}^{t}\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}|^{k}]\big{)}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\leq\frac{1}{2}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}+\frac{\tilde{c}(j)^{2}}{2}\int_{0}^{t}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\text{\rm{d}}s,$
for some $\tilde{c}(j)\geq c(j)$ as in (3.31), by Gronwall’s inequality we find a constant $c_{5}(\mu,\phi)>0$ such that
(3.48) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{k}\Big{]}\leq c_{5}(\mu,\phi),\ \ \varepsilon\in(0,1],$
so that (3.47) implies
(3.49) $\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\tilde{v}_{s}^{\varepsilon}|^{j}\Big{|}\mathscr{F}_{0}\Big{]}\leq c_{6}(\mu,\phi)+c(j)|\phi(X_{0}^{\mu})|^{j},\ \ \varepsilon\in(0,1]$
for some constant $c_{6}(\mu,\phi)>0.$
(d) For $|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}$. Now, let
$\psi_{t}^{\varepsilon}:=\,$ $\mathbb{E}\Big{[}\big{\langle}\{D^{L}B_{t}(z,\cdot,P_{t}V(\mu))\}(P_{t}^{*}\mu)(X_{t}^{\mu}),(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}^{\varepsilon}\big{\rangle}\Big{]}|_{z=X_{t}^{\mu}}$
+$\qquad+\big{\{}\nabla^{\mathbb{B}}_{v_{t}^{\varepsilon}}B_{t}(X_{t}^{\mu},P_{t}^{*}\mu,\cdot)\big{\}}(P_{t}V(\mu)).$
Then (3.46) can be reformulated as
$\tilde{v}_{t}^{\varepsilon}-\tilde{v}_{0}^{\varepsilon}$
$=\int_{0}^{t}\Big{\{}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})+\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu})\Big{\}}\text{\rm{d}}s$
$\quad+\int_{0}^{t}\big{[}\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})\big{]}\text{\rm{d}}W_{s}+\alpha_{t}^{\varepsilon},\ \ t\in[0,T],$
where for $t\in[0,T],$
$\alpha_{t}^{\varepsilon}:=\int_{0}^{t}\xi_{s}^{\varepsilon}\text{\rm{d}}s+\int_{0}^{t}\tilde{\xi}_{s}^{\varepsilon}\,\text{\rm{d}}W_{s},$
$\xi_{s}^{\varepsilon}:=\frac{1}{\varepsilon}\{[B_{s}(\cdot,P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))+\lambda_{0}u_{s}](X_{s}^{\mu_{\varepsilon}})-[B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}](X_{s}^{\mu})\}$
$\qquad\qquad+\nabla_{\frac{1}{\varepsilon}\{B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon}},P_{s}^{*}\mu,P_{s}V(\mu))\}}u_{s}(X_{s}^{\mu_{\varepsilon}})$
$\qquad-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\big{[}B_{s}(\cdot,P_{s}^{*}\mu,P_{s}V(\mu))+\lambda_{0}u_{s}\big{]}(X_{s}^{\mu})-\nabla_{\psi_{s}^{\varepsilon}}\Theta_{s}(X_{s}^{\mu}),$
$\tilde{\xi}_{s}^{\varepsilon}:=\frac{\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu_{\varepsilon}})-\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu})}{\varepsilon}-\nabla_{(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\tilde{v}_{s}^{\varepsilon}}\{(\nabla\Theta_{s})\sigma_{s}\}(X_{s}^{\mu}).$
By the same argument leading to [29, (4.23),(4.26)] and (3.30) above, (B) together with (2.9), (3.45), (3.48), (3.49) implies
$\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,T]}|\alpha_{s}(\varepsilon)|^{j}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ j\geq 1,$
and by Jensen’s inequality we find a constant $c>0$ such that
$\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\leq c|\tilde{v}_{0}^{\varepsilon}-\tilde{v}_{0}|^{k}+c\bigg{(}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\alpha_{s}(\varepsilon)|^{2k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{)}^{\frac{1}{2}}$
(3.50) $\quad+c\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\bigg{(}\int_{0}^{t}\Big{\{}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in[0,T].$
(e) For $\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$. Similarly to (d), for $N_{s}$ and $\tilde{N}_{s}$ in (3.13), we have
(3.51) $\begin{split}&\frac{R_{t}^{\varepsilon}-1}{\varepsilon}=\int_{0}^{t}R_{s}^{\varepsilon}\big{\langle}\varepsilon^{-1}\eta_{s}^{\varepsilon},\text{\rm{d}}W_{s}\big{\rangle}\\ &=\int_{0}^{t}R_{s}^{\varepsilon}\Big{\langle}\frac{\zeta_{s}(X_{s}^{\mu_{\varepsilon},\mu})[B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))]}{\varepsilon},\ \text{\rm{d}}W_{s}\Big{\rangle}\\ &=h_{t}(\varepsilon)+\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon})+\tilde{N}_{s}(\tilde{v}^{\varepsilon})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle},\end{split}$
where
$h_{t}(\varepsilon):=\int_{0}^{t}\Big{\langle}$ $\zeta_{s}(X_{s}^{\mu})\Big{\{}R_{s}^{\varepsilon}\frac{B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu_{\varepsilon},P_{s}V(\mu_{\varepsilon}))-B_{s}(X_{s}^{\mu_{\varepsilon},\mu},P_{s}^{*}\mu,P_{s}V(\mu))}{\varepsilon}$
$\qquad\qquad-N_{s}(v^{\varepsilon})-\tilde{N}_{s}(\tilde{v}^{\varepsilon})\Big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}$
satisfies
(3.52) $\lim_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{t\in[0,T]}|h_{t}(\varepsilon)|^{p}\Big{|}\mathscr{F}_{0}\Big{]}=0,\ \ p\geq 1.$
Noting that $X_{t}^{\mu}=X_{t}^{\mu,\mu}$, by (3.25) we find a constant $c(\mu,\phi)>0$ such that
(3.53) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|X_{t}^{\mu_{\varepsilon},\mu}-X_{t}^{\mu}|^{k}\Big{]}\leq c(\mu,\phi)\varepsilon^{k},\ \ \varepsilon\in[0,1],\mu\in\mathscr{P}_{k}.$
Combining these with (3.42), (3.45) and (3.44), we find a constant $k_{1}(\mu,\phi)>0$ such that
(3.54) $\Big{\|}v_{t}^{\varepsilon,2}-\frac{1}{\varepsilon}\mathbb{E}[V(X_{t}^{\mu})(R_{t}^{\varepsilon}-1)]\Big{\|}_{\mathbb{B}}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1).$
By (3.14) for $v_{t}=v_{t}(\mu,\phi),\tilde{v}_{t}=\tilde{v}_{t}(\mu,\phi)$, and by the definition of $\hat{v}_{t}$ in (3.35), we have
$\hat{v}_{t}$ $=I_{t}^{V}(\mu,\phi)+\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(\hat{v})+\tilde{N}_{s}(\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}.$
Combining this with (3.33), (3.38), (3.52), (3.51), (3.54) and the same argument to (3.27), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times[0,1]\rightarrow(0,\infty)$ with
(3.55) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0,$
@@
$\|v_{t}^{\varepsilon}-\hat{v}_{t}\|_{\mathbb{B}}$
(3.56) $\leq\tilde{h}_{t}(\varepsilon)+\bigg{\|}\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}(v^{\varepsilon}-\hat{v})+\tilde{N}_{s}(\tilde{v}^{\varepsilon}-\tilde{v})\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\mathbb{B}}$
$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\Big{\{}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}+\big{(}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\big{)}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
(f) Proof of (3.35). By (3.49)-(3.3) and Fatou’s lemma, we see that
$\begin{split}&\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{]},\ \ \tilde{\beta}_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}\end{split}$
satisfy
(3.57) $\begin{split}&\beta_{t}\leq\mathbb{E}\bigg{[}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}\Big{[}\sup_{s\in[0,t]}|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}\Big{|}\mathscr{F}_{0}\Big{]}\bigg{]}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\limsup_{\varepsilon\rightarrow 0}\frac{s}{\alpha(s^{\frac{1}{2}})^{2}}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}\\ &\qquad\quad+c\bigg{(}\int_{0}^{t}\Big{\{}\limsup_{\varepsilon\rightarrow 0}\mathbb{E}[|\tilde{v}_{s}^{\varepsilon}-\tilde{v}_{s}|^{k}]\Big{\}}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\\ &\leq c\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\,\tilde{\beta}_{s}^{2}\text{\rm{d}}s\bigg{)}^{\frac{k}{2}}+c\left(\int_{0}^{t}\beta_{s}^{2}\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ t\in[0,T].\end{split}$
Noting that (3.16), (3.34) and (3.45) imply that $\tilde{\beta}_{t}$ satisfies
$\sup_{t\in(0,T]}\tilde{\beta}_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon}-\hat{v}_{s}\|_{\mathbb{B}}=:\tilde{c}(\mu,\phi)<\infty,$
@@
$\tilde{\beta}_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\Big{\{}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\tilde{\beta}_{s}^{2}+\beta_{s}^{\frac{2}{k}}\Big{\}}\text{\rm{d}}s,\ \ t\in(0,T],$
@@
Combining this with (3.57), $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.35) and hence finish the proof. ∎

### v4

@@ 2. $(2)$
(3.49) ×$\qquad\qquad\quad\times\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})\big{\{}[v^{\varepsilon,1}_{s}-\hat{v}_{s}^{1}](X_{s}^{x,\mu},P_{s}^{\ast}\mu)\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\infty}\mu(\text{\rm{d}}x)=0.$
Moreover, combining (3.3) with $[U_{t}(y,\nu,\cdot)]_{\alpha}\leq 1,$ and (2.9) for $p=k^{\ast}$, we obtain
$\Big{\|}v_{t}^{\varepsilon,2}-\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x)\Big{\|}_{\infty}$
$\leq\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\mathbb{E}[\alpha(|X_{t}^{x,\mu}-X_{t}^{x+\varepsilon\phi(x),\mu}|)|R_{t}^{\varepsilon,x}-1|]\mu(\text{\rm{d}}x)$
$\leq\frac{1}{\varepsilon}\int_{\mathbb{R}^{d}}\left\{\Big{(}\mathbb{E}[|R_{t}^{\varepsilon,x}-1|^{k^{*}}]\Big{)}^{\frac{1}{k^{*}}}\alpha\Big{(}\big{(}\mathbb{E}[|X_{t}^{x,\mu}-X_{t}^{x+\varepsilon\phi(x),\mu}|^{k}]\big{)}^{\frac{1}{k}}\Big{)}\right\}\mu(\text{\rm{d}}x).$
This together with (3.45) and (3.39) yields that for some constant $k_{1}(\mu,\phi)>0$,
$\Big{\|}v_{t}^{\varepsilon,2}-\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}\mathbb{E}[U_{t}(\cdot,\cdot,X_{t}^{x,\mu})(R_{t}^{\varepsilon,x}-1)]\mu(\text{\rm{d}}x)\Big{\|}_{\infty}\leq k_{1}(\mu,\phi){{\alpha(\varepsilon)}},\ \ t\in[0,T],\ \varepsilon\in(0,1].$
Combining this with (3.30), (3.3), (3.43), (3.3), (3.42), (3.48), and the argument leading to (3.22), we find a constant $k_{2}(\mu,\phi)$ and a measurable function $\tilde{h}:(0,T]\times(0,1]\rightarrow(0,\infty)$ with
(3.50) $\sup_{\varepsilon\in(0,1],t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)\leq k_{2}(\mu,\phi),\ \ \lim_{\varepsilon\rightarrow 0}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\tilde{h}_{t}(\varepsilon)=0$
@@
$\left\|v_{t}^{\varepsilon,2}-\hat{v}_{t}^{2}\right\|_{\infty}$ $\leq\tilde{h}_{t}(\varepsilon)+\int_{\mathbb{R}^{d}}\bigg{\|}\mathbb{E}\bigg{[}U_{t}(\cdot,\cdot,X_{t}^{x,\mu})$
(3.51) ×$\qquad\qquad\quad\times\int_{0}^{t}\Big{\langle}\{\zeta_{s}[v^{\varepsilon,2}_{s}-\hat{v}_{s}^{2}]\}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\bigg{\|}_{\infty}\mu(\text{\rm{d}}x)$
$\leq\tilde{h}_{t}(\varepsilon)+k_{2}(\mu,\phi)\bigg{(}\int_{0}^{t}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}^{2}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ t\in(0,T].$
(d) Proof of (3.31). Let
$\begin{split}\beta_{t}:=\limsup_{\varepsilon\rightarrow 0}\sup_{s\in(0,t]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}.\end{split}$
Noting that (3.18), (3.26), (3.27), (3.21) and (3.40) imply that $\beta_{t}$ satisfies
$\sup_{t\in(0,T]}\beta_{t}\leq\sup_{\varepsilon\in(0,1]}\sup_{s\in(0,T]}\frac{\sqrt{s}}{\alpha(s^{\frac{1}{2}})}\|v_{s}^{\varepsilon,2}-\hat{v}_{s}^{2}\|_{\infty}=:\tilde{c}(\mu,\phi)<\infty,$
@@
$\beta_{t}^{2}\leq Ck_{2}(\mu,\phi)^{2}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s}\beta_{s}^{2}\text{\rm{d}}s,\ \ t\in(0,T],$
@@
Combining this with $\int_{0}^{T}\frac{\alpha(t^{\frac{1}{2}})^{2}}{t}\text{\rm{d}}t<\infty$, and applying Gronwall’s inequality, we prove (3.31), which together with (3.34) completes the proof. ∎